OpenStudy (anonymous):

use a calculator to find the value of sec-1(6.2) how do i enter this in my ti83

OpenStudy (anonymous):

you are trying to solve \[\frac{1}{\cos(x)}=6.2\] for x so \[\cos(x)=\frac{1}{6.2}\] and therefore \[x=\cos^{-1}(\frac{1}{6.2})\]

myininaya (myininaya):

 \[\cos(\theta)=\frac{adj}{hyp} ; \sec(\theta)=\frac{hyp}{adj}\] remember \[\frac{1}{\cos(\theta)}=\sec(\theta)\] \[\theta=\sec^{-1}(\frac{1}{\cos(\theta)})=\sec^{-1}(\frac{hyp}{adj})\] \[\cos(\theta)=\frac{1}{\sec(\theta)}\] \[\theta=\cos^{-1}(\frac{1}{\sec(\theta)})=\cos^{-1}(\frac{adj}{hyp})\] => \[\cos^{-1}(\frac{adj}{hyp})=\sec^{-1}(\frac{hyp}{adj})\]